OpenStudy (anonymous):

What is the pressure in atmospheres of 512g of N2 gas in an 12.00L container at 37.0oC ?

OpenStudy (abmon98):

Use ideal gas equation PV=nRT

OpenStudy (anonymous):

p*12=(18.27)*(.297)*(37)?

OpenStudy (abmon98):

Change 37 degree celsium to Kelvin by adding to 273 (273+37 C)=310 Kelvin R-gas constant-0.08205746(14 L atm K−1 mol Number of moles=Mass(g)/Molar Mass(g/mol) N of N2=512/28=18.28 moles P is measured in atm P=18.28*0.0820*310/12